Problem 10-8AB Effective Interest: Amortization of bond discount LO P5 Legacy issues $710,000 of 8.0%, four-year bonds datedRequired 1 Required 2 Required 3 Required 4 Complete the below table to calculate the total bond interest expense to be recogRequired 1 Required 2 Required 3 Required 4 Prepare an effective interest amortization table for the bonds first two years.Required 1 Required 2 Required 3 Required 4 Prepare the journal entries to record the first two interest payments. View transJournal entry worksheet < 1 2 Record the second interest payment on December 31. Note: Enter debits before credits. General J
Answer #1

Solution

Legacy

  1. Entry to record bond’s issuance:

Date

Account Titles and Explanation

Debit

Credit

Jan 1, 2019

Cash

$621,812

Discount on Bonds Payable

$88,188

Bonds Payable

$710,000

(To record bond issuance and discount)

Computations:

  1. Discount on issue of bonds = face value – issue price

= 710,000 – 621,812 = $88,188

  1. Calculation of bond interest expense:

Semi-annual interest expense = 710,000 x 8% x 6/12 = $28,400

Amount Repaid:

8

payments of

$28,400

$227,200

par value at maturity

$710,000

total repaid

$937,200

Less: amount borrowed

$621,812

Total bond interest expense

$315,388

  1. Effective interest amortization table for the bond’s first two years:

Date

Cash Paid

Interest Expense

Discount Amortization

Carrying Value of Bond

Jan 1, 2019

-

-

-

$621,812

June 30, 2019

$28,400

$37,308

$8,909

$630,721

Dec 31, 2019

$28,400

$37,843

$9,443

$640,164

June 30, 2020

$28,400

$38,410

$10,010

$650,174

Dec 31, 2020

$28,400

$39,010

$10,610

$660,784

June 30, 2021

$28,400

$39,647

$11,247

$672,031

Dec 31, 2021

$28,400

$40,322

$11,922

$683,953

June 30, 2022

$28,400

$41,037

$12,637

$696,590

Dec 31, 2022

$28,400

$41,795

$13,395

$709,985

Computations:

Interest expense for each semi-annual period is computed as follows,

Interest expense = carrying value of bond x market rate of interest x 6/12

June 30, 2019 = 621,812 x 12% x 6/12 = $37,308

Discount amortization = interest expense – cash paid = 37,308 – 28,400 = $8,909

Carrying value of bond = beg. Balance + discount amortized

= 621,812 + 8,909 = $630,721

Total interest expense = $315,372

Total cash paid = $227,200

Total discount amortization = $88,173

  1. Entries to record the first two interest payments:

Date

Account Titles and Explanation

Debit

Credit

June 30, 2019

Interest Expense

$37,308

Discount on Bonds Payable

$8,908

Cash

$28,400

(To record first interest payment)

Dec 31, 2019

Interest Expense

$37,843

Discount on Bonds Payable

$9,443

Cash

$28,400

(To record second interest payment)